Petra Paddles Company is budgeting for Quarter 3 (July, August and September). These are busy months, so it is important to get the numbers right and to come up with a good plan. This company sells kayaks and does 4-hour Kayak tours. Petra charges $120 for tours and kayaks are sold as a package for $3,000 each (includes a life jacket, a paddle, safety gear and the kayak). Based on last year’s sales, Petra estimates that she will sell 12 kayaks in July, 7 in August and 5 in September. All of these are expected to be cash sales. Each kayak package costs $2,700. 

Every year, a large group of 20 people returns for a tour in July. This group has already prepaid for their booking. This tour and all other projected bookings are included below:

  

July

August

September

Budgeted number of tours to complete

445+20 =465

400

250


I


1
Expert's answer
2021-04-15T10:53:09-0400

Budgeting

July

"Total=(12 \times2,700)+(\frac{465}{1110}\times54000)=55,062"

August

"Total=(7\times2,700)+(\frac{400}{1110}\times54000)=38,359"

September

"Total=(5\times2,700)+(\frac{250}{1110}\times54000)=25,662"


Working

"Total revenue=20\times2,700=54,000"